INFRARED-ULTRAVIOLET DOUBLE RESONANCE SPECTROSCOPY OF ACETYLENE: UNRAVEL-LING THE MYSTERIES IN THE ν_{CC} + 3ν_{CH} REGION AT 11 600 cm^{−1}USING THE CLUSTER/POLYAD MODEL

\underline{ANGELA P. MILCE} and BRIAN J. ORR, School of Chemistry and Centre for Lasers and Applications, Macquarie University, Sydney, NSW, Australia 2109.

Time-resolved fluorescence-detected infrared-ultraviolet optical double resonance (IR-UV DR) experiments have been used to obtain detailed measurements of spectroscopic and dynamical processes in the ”ν_{CC} +3ν_{CH}” region of acetylene (C_{2}H_{2}) at 11 600 cm^{−1}. The rotational energy transfer (RET) and vibrational energy transfer (V-V) kinetic data collected have already been well fit to a detailed rate-equation model that incorporates empirical exponential-gap fitting laws. ^{a}

In order to explain some of the unusual effects observed in this region we now use the cluster/polyad model that has already successfully interpreted the vibrational energy pattern in C_{2}H_{2} up to 12 000 cm^{−1}. ^{b} Our work emphasises rotational perturbations, after including the rotational l-type resonance interaction. This enables us to identify the intramolecular couplings responsible for enhancement of rotationally-resolved V-V energy transfer between coupled rovibrational levels of the ”ν_{CC}+ 3ν_{CH}” manifold. The model also provides qualitative evidence that Coriolis-type perturbations are responsible for other unusual symmetry-breaking phenomena in this region.

aA. P. Milce, H. -D. Barth and B. J. Orr, J. Chem. Phys. \underline{100}, 2398 (1994); A. P. Milce and B. J. Orr, J. Chem. Phys. \underline{104}, 6423 (1996); ibid. \underline{106}, (1997) in print.

bM. A. Temsamani and M. Herman, J. Chem. Phys. \underline{102}, 6371 (1995); ibid. \underline{105}, 1355 (1996); M. A. Temsamani, M. Herman, S. A. B. Solina, J. P. O’Brien and R. W. Field, J. Chem. Phys. \underline{105}, 11357 (1996).